OpenStudy (anonymous):

(5+2i)(4+3i) Can someone please explain me in steps how to simplify it?

OpenStudy (anonymous):

you have a choice here you can either just multiply out like in "foil" (ugh) or you can use a formula for multiplying complex numbers. which do you choose (or both)?

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\] you should really learn this if you are going to see these more often.

OpenStudy (anonymous):

The one you may think is easier, it seems the foil method isn't your choice anyway. lol

OpenStudy (anonymous):

so in your case you will get \[(5+2i)(4+3i)=(5\times 4-2\times 3)+(5\times 3+2\times 4)i\] \[(20-6)+(15+8)i\] \[14+23i\]

OpenStudy (anonymous):

(5+2i)(4+3i) (2i)(3i)=6i^2=-6 (2i)(4)=8i (3i)(5)=15i (4)(5)=20 14+23i

OpenStudy (anonymous):

if this is too hard to remember multiply out like you would a two term by two term polynomial, and then at the end recall that \[i^2=-1\] so you can put \[5\times 4+5\times 3i+2\times 4i+6\times i^2\] \[20+15i+8i-6\] \[14+23i\] it is your choice

OpenStudy (anonymous):

Thanks that formula was really helpful. :)